# MuSACo: Multimodal Subject-Specific Selection and Adaptation for Expression Recognition with Co-Training

### v1

@@ 6.4 Datasets
Biovid Heat Pain [38] dataset consists of 87 subjects captured in a controlled environment. It consists of five classes, including: No-Pain, PA-1, PA-2, PA-3, and PA-4. Every individual recorded 20 videos per class, corresponding to 100 videos per subject. We follow the same protocol as [40], where it eliminates the initial 2 seconds from the video, which does not show any spike indicating pain. In our experiment, we follow [45] to consider 77 subjects as sources and 10 subjects as targets: Sub-1 (081014_w_27), Sub-2 (101609_m_36), Sub-3 (112009_w_43), Sub-4 (091809_w_43), Sub-5 (071309_w_21), Sub-6 (073114_m_25), Sub-7 (080314_w_25), Sub-8 (073109_w_28), Sub-9 (100909_w_65), Sub-10 (081609_w_40). StressID [4] dataset comprised 65 (47 men and 18 women) subjects taken in a lab-controlled environment. Each participant was exposed to 11 tasks, grouped into 4 categories: watching emotional videos, breathing, interactive tasks, and relaxation. These tasks were captured using three different modalities (visual, physiological, and audio). Due to the unavailability of certain modalities in 11 of the participants, we only consider 54 individuals, which includes all the modalities. In our experiments, we take 44 subjects as sources and 10 subjects as targets. The target subject selection follows a 70% (men) and 30% (women) ratio, reflecting the higher number of male participants in the dataset. For each participant, there are up to 11 task-specific videos, from which we extracted frames at 1 fps for per-image expression classification. Tab. 4 shows the selected target subject demographics (ID, Gender) and the total number of samples.
@@
6.5 Ablation
(a) Impact of Weighting Hyperparameters. We performed a sequential weight sensitivity analysis for the key loss terms, $\mathcal{L}_{unsup}^{t}$, $\mathcal{L}_{aw}$, and $\mathcal{L}_{agn}$ shown in Tab. 5). First, we varied the weight of the target PL loss while keeping the others fixed at 1. Using the best result, we tuned the class-aware alignment loss, followed by the class-agnostic loss, each time fixing the previously selected best weights. This approach highlights the individual contribution of each loss to overall performance.
Weight Accuracy
@@ 6.5 Ablation
(b) Ablation on Target PLs Threshold ($\tau_{pl}$). We conducted an ablation to select a target PLs threshold ($\tau_{pl}$) value as shown in the Tab. 7. We found that $\tau_{pl}$ = 0.95 produced the best trade-off, balancing label quality with sample quantity. To mitigate the issue of excluding useful but uncertain samples when using a high threshold, we introduce a domain-agnostic loss that ensures the model will also learn from the non-confident samples. Note that target ground-truth labels were used for this analysis to evaluate the effectiveness of pseudo-labels.
@@
(c) Source Subject Selection Analysis Tab. 6 shows the number of selected source subjects for each of the 10 target subjects for the Biovid dataset for various threshold ($\tau_{ss}$) values. As $\tau_{ss}$ increases, fewer source subjects are selected, reducing potential noise from irrelevant domains. Notably, the accuracy improves when transitioning from $\tau=0$ to moderate thresholds (e.g., $\tau_{ss}=0.45$ or $0.55$), demonstrating that selective inclusion of source data benefits generalization. The best average accuracy of 39.89% is achieved with $\tau_{ss}=0.55$, using approximately 21 subjects on average.
@@
Table 8: Ablation on disentanglement module for visual and physiological modalities for Biovid dataset.
@@
CAN 59 19,625 8.03 34.6
Sub-based 59 234 7.36 34.6
CMSDA 271 4.07 36.8
@@
6.6 MuSACo

### v2

@@ 9.2 StressID
[chaptoukaev2023stressid] dataset comprised 65 (47 men and 18 women) subjects taken in a lab-controlled environment. Each participant was exposed to 11 tasks, grouped into 4 categories: watching emotional videos, breathing, interactive tasks, and relaxation. These tasks were captured using three different modalities (visual, physiological, and audio). Due to the unavailability of certain modalities in 11 of the participants, we only consider 54 individuals, which includes all the modalities. In our experiments, we take 44 subjects as sources and 10 subjects as targets. The target subject selection follows a 70% (men) and 30% (women) ratio, reflecting the higher number of male participants in the dataset. For each participant, there are up to 11 task-specific videos, from which we extracted frames at 1 fps for per-image expression classification. Tab. 6 shows the selected target subject demographics (ID, Gender) and the total number of samples. Tab. 7 shows the label distribution of StressID datasets. It exhibits a noticeable imbalance between stress and no-stress samples across subjects. Only two subjects, Sub-1 and Sub-8, are somewhat balanced, while others are skewed.
9.3 Behavioural Ambivalence/Hesitancy (BAH)
The Behavioural Ambivalence/Hesitancy (BAH) [gonzalez2025bah] dataset is the first multimodal resource designed for subject-specific recognition of ambivalence and hesitancy (A/H) in real-world settings. It consists of 1,118 videos (8.26 hours, including 1.5 hours of A/H) from 224 participants across 9 Canadian provinces, covering diverse ages and ethnicities. Participants answered 7 questions, some explicitly eliciting A/H, while being recorded via webcam and microphone, providing multimodal signals such as facial expressions, vocal cues, and transcripts. Expert annotations mark A/H occurrences at both frame- and video-level, with aligned face crops, transcripts, and participant metadata also included. Unlike controlled lab datasets (e.g., BioVid, StressID), BAH captures natural, in-the-wild behaviours, making it uniquely challenging and valuable for multimodal domain adaptation. In our experiments, we use 143 training data from BAH as subjects to consider as source domains and use 5 subjects from the test-set of BAH as targets: Sub-1 (82711), Sub-2 (82585), Sub-3 (82683), Sub-4 (82708), and Sub-5 (82632), enabling a realistic evaluation of MuSACo under noisy, heterogeneous conditions. The label distribution statistics for the test set of Stress-ID were shown in Tab. 8 It also shows a significant imbalance in ambivalence vs. non-ambivalence across most subjects, with only Sub-1 (82711) and Sub-8 (82683) being closer to balanced.
Subjects ID NA PA1 PA2 PA3 PA4 Total
Sub-1 081014_w_27 287 285 311 305 312 1500
Sub-2 101609_m_36 325 278 323 274 300 1500
Sub-3 112009_w_43 304 282 296 310 308 1500
Sub-4 091809_w_43 265 316 295 298 302 1476
Sub-5 071309_w_21 299 297 295 304 297 1492
Sub-6 073114_m_25 295 310 320 283 292 1500
Sub-7 080314_w_25 311 318 288 276 307 1500
Sub-8 073109_w_28 314 300 300 272 303 1489
Sub-9 100909_w_65 298 298 313 279 312 1500
Sub-10 081609_w_40 300 297 313 286 303 1499
Table 5: Label distribution for BioVid dataset across target subjects test-set.
@@ 9.3 Behavioural Ambivalence/Hesitancy (BAH)
Subjects ID No-Stress Stress Total
Sub-1 kycf 118 86 204
Sub-2 uymz 148 28 176
Sub-3 h8s1 128 83 211
Sub-4 ctzy 60 148 208
Sub-5 p9i3 125 83 208
Sub-6 7h5u 127 57 184
Sub-7 g7r2 175 33 208
Sub-8 b9w0 105 100 205
Sub-9 r3zm 63 120 183
Sub-10 x1q3 59 142 201
Table 7: Label distribution for StressID dataset across target subjects test-set.
Subjects ID No-Amb Amb Total
Sub-1 82711 433 874 1307
Sub-2 82687 319 207 526
Sub-3 82585 575 85 660
Sub-4 82592 737 93 830
Sub-5 82598 1199 189 1388
Sub-6 82632 854 243 1097
Sub-7 82681 354 207 561
Sub-8 82683 445 456 901
Sub-9 82708 249 104 353
Sub-10 82714 231 234 465
Table 8: Label distribution for BAH dataset across target subjects test-set.
10 Ablation
10.1 Impact of Weighting Hyperparameters.
We performed a sequential weight sensitivity analysis for the key loss terms, $\mathcal{L}_{unsup}^{t}$, $\mathcal{L}_{aw}$, and $\mathcal{L}_{agn}$ shown in Tab. 9). The weight sensitivity analysis was performed on three representative subjects from the BioVid dataset: a young woman (081014_w_27), a young man (073114_m_25), and an older woman (100909_w_65). The reported results were the average of all three subjects, and each experiment was run for 10 epochs. These subjects were chosen to reflect the diversity of the dataset across age and gender groups. First, we varied the weight of the target PL loss while keeping the others fixed at 1. Using the best result, we tuned the class-aware alignment loss, followed by the class-agnostic loss, each time fixing the previously selected best weights. This sequential tuning strategy allowed us to identify the most effective configuration while minimizing computation cost. The best configuration obtained was $\gamma_{t}=1$, $\alpha=0.1$, and $\beta=0.5$. All experiments in the paper were conducted using this configuration to ensure consistency across results.
Weights Accuracy
@@ 10 Ablation
10.2 Source Subject Selection Analysis
Tab. 10 shows the number of selected source subjects for each of the 10 target subjects for the Biovid dataset for various threshold ($\tau_{ss}$) values. As $\tau_{ss}$ increases, fewer source subjects are selected, reducing potential noise from irrelevant domains. Notably, the accuracy improves when transitioning from $\tau=0$ to moderate thresholds (e.g., $\tau_{ss}=0.45$ or $0.55$), demonstrating that selective inclusion of source data benefits generalization. The best average accuracy of 39.89% is achieved with $\tau_{ss}=0.55$, using approximately 21 subjects on average. To assess the robustness of the source subject selection threshold ($\tau_{ss}$), we conducted experiments on the Behavioural Ambivalence/Hesitancy (BAH) [gonzalez2025bah] dataset, which differs substantially from lab-controlled datasets (BioVid and Stress) as it consists of self-recorded, in-the-wild videos. Table 11 shows results across thresholds from 0.45 to 0.65. The analysis reveals that accuracy remains stable in the 0.45–0.55 range, with the best performance achieved at $\tau_{ss}=0.55$. Higher thresholds reduce the number of selected sources and lead to performance degradation due to over-pruning. These findings confirm that $\tau_{ss}$ is not highly sensitive to fine-tuning and that a default value around 0.55 provides a reliable trade-off between accuracy and source diversity across datasets.
10.3 Ablation on Target PLs Threshold ($\tau_{pl}$).
Tab. 12, and found that $\tau_{0}=0.95$ provided the best trade-off between label quality and sample quantity. Rather than fixing this value, we employ a confidence-annealing schedule, where $\tau_{pl}$ is decreased by $0.01$ every $N$ epochs, allowing more target samples to be included as the model stabilizes. This ensures early training benefits from high-confidence labels, while later stages incorporate a broader set of samples. To further mitigate the exclusion of useful but uncertain samples, we introduce a domain-agnostic loss that allows learning from low-confidence examples as well. Target ground-truth labels were only used here to evaluate pseudo-label quality.
$\tau_{ss}$ Sub-1 Sub-2 Sub-3 Sub-4 Sub-5 Avg SS Avg Acc
0.45 44 33 39 41 39 39.2 67.1
0.50 42 31 36 40 36 37.0 68.2
0.55 41 30 35 38 34 35.6 68.9
0.60 39 28 33 37 33 34.0 66.6
0.65 38 27 30 36 30 32.2 62.5
Table 11: Impact of varying source subject selection threshold $\tau_{ss}$ on the BAH dataset.
@@ 10.2 Source Subject Selection Analysis
Table 13: Ablation on the disentanglement module for visual and physiological modalities for the Biovid dataset.
10.4 Impact of Disentanglement
conducted separately for visual and physiological modalities on the BioVid dataset, using 77 source and 10 target subjects. Results (Fig. 6, Tab. 13) show that incorporating the KNIFE-based disentanglement improves performance for both modalities. Without disentanglement, identity information interferes with expression classification, resulting in reduced accuracy. Adding the KNIFE-loss mitigates this, and introducing an identity head yields further improvements, confirming that explicitly removing identity bias enhances modality-specific expression features.
10.5 Computational Complexity
@@
CAN [kang2020contrastive] 59 19,625 8.03 34.6
Sub-based [zeeshan2024subject] 59 234 7.36 34.6
CMSDA [scalbert2021multi] 271 4.07 36.8
@@
$\mathcal{L}^{s}$ $\mathcal{L}^{t}$ $\mathcal{L}_{aw}$ $\mathcal{L}_{agn}$ Avg Acc
32.1
36.5
38.0
37.2
40.7
Table 15: Impact of individual loss components under equally weighted hyperparameter.
Settings Methods Target Subjects
Sub-1 Sub-2 Sub-3 Sub-4 Sub-5 Sub-6 Sub-7 Sub-8 Sub-9 Sub-10 Avg F1
Lower Bound
Visual-only
Physio-only
Fusion
63.1
36.6
57.3
24.6
45.6
32.1
61.2
37.7
50.0
37.8
22.3
36.6
58.3
37.5
51.5
50.8
40.8
63.4
43.2
45.6
68.4
36.6
33.8
39.5
46.9
25.6
52.6
54.8
22.7
53.4
47.7
34.8
50.5
$M$-UDA
DANN [ganin2016domain]
CDAN [long2018conditional]
MMD [sejdinovic2013equivalence]
MuSACo (UDA)
61.2
57.1
57.1
37.6
45.3
43.2
54.0
49.3
55.6
64.3
61.3
74.1
51.3
46.6
40.3
86.2
53.6
48.3
45.6
55.7
56.6
58.3
55.6
61.5
56.6
59.6
55.6
64.8
47.6
45.6
55.6
53.4
62.3
62.2
55.4
74.4
38.6
39.6
49.3
50.0
52.9
52.5
53.0
60.7
$M$-MSDA
CAN [kang2020contrastive]
Sub-basedtop-k [zeeshan2024subject]
MuSACo (MSDA)
65.3
68.9
76.1
45.4
45.3
59.6
54.6
74.2
69.3
42.3
46.6
53.3
51.2
55.2
58.6
58.6
62.6
67.2
66.6
56.6
75.2
54.2
39.3
48.3
59.3
52.2
74.6
34.6
39.9
40.2
53.2
54.1
62.2
Upper Bound Fine-tuning 86.3 43.8 94.4 97.7 79.6 75.6 45.6 93.0 91.0 97.0 80.4
Table 16: Subject-wise F1 scores of StressID dataset across 10 target subjects.
Settings Methods Target Subjects
Sub-1 Sub-2 Sub-3 Sub-4 Sub-5 Sub-6 Sub-7 Sub-8 Sub-9 Sub-10 Avg F1
Lower Bound
Visual-only
Physio-only
Fusion
29.8
38.1
41.2
26.1
32.5
27.2
9.1
39.1
24.9
28.1
15.4
32.2
19.3
10.4
29.5
21.6
29.0
25.8
25.7
14.3
34.7
15.3
34.6
24.3
16.5
28.7
22.3
16.1
22.0
18.9
20.8
26.4
28.1
$M$-UDA
DANN [ganin2016domain]
CDAN [long2018conditional]
MMD [sejdinovic2013equivalence]
MuSACo (UDA)
32.5
28.5
28.3
35.6
25.6
24.3
20.7
21.7
29.5
25.5
29.6
33.3
33.6
29.2
32.4
36.2
28.3
25.2
28.4
28.7
25.1
24.2
29.5
16.0
29.6
28.6
32.1
31.2
24.2
23.4
28.6
29.4
25.2
23.1
25.4
35.4
20.2
20.5
26.5
17.6
27.4
25.3
28.2
28.5
$M$-MSDA
CAN [kang2020contrastive]
Sub-basedtop-k [zeeshan2024subject]
CMSDA [scalbert2021multi]
MuSACo (MSDA)
34.6
34.6
31.2
41.6
30.3
37.6
32.5
33.2
22.3
22.4
21.2
49.6
24.6
20.5
31.2
27.8
27.9
34.5
32.3
25.6
26.3
24.6
30.1
36.5
28.6
36.8
30.2
29.4
26.6
24.3
34.1
35.4
33.2
30.6
38.9
49.0
18.1
17.6
18.5
29.2
27.3
28.4
30.0
35.7
Upper Bound Fine-tuning 81.2 74.3 67.6 68.2 70.3 61.4 74.2 70.6 74.0 49.5 69.1
Table 17: Subject-wise F1 scores of BioVid dataset on 10 target subjects.
Settings Methods Sub-1 Sub-2 Sub-3 Sub-4 Sub-5 Avg F1
Lower Bound Visual-only 25.2 46.2 35.2 41.3 43.2 38.2
Audio-only 27.4 34.4 34.1 40.3 44.1 36.1
Fusion 26.3 44.5 34.7 41.9 44.6 38.4
$M$-UDA MMD [sejdinovic2013equivalence] 28.2 44.5 39.3 38.7 43.2 38.8
MuSACo (UDA) 26.1 46.1 35.6 42.2 43.5 38.7
$M$-MSDA Sub-basedtop-k [zeeshan2024subject] 24.6 45.4 36.8 42.7 44.9 38.9
MuSACo (MSDA) 28.2 48.6 39.4 44.2 45.4 41.2
Table 18: Subject-wise F1 scores on the BAH dataset across five target subjects. MuSACo MSDA consistently outperforms baselines and other UDA/MSDA methods.
10.6 Impact of Difference Loss Components with Equal Weights
In addition to the tuned-weight ablation reported in the main paper, we also conducted a diagnostic experiment where all loss weights were fixed to 1. This isolates the relative impact of each component without any hyperparameter tuning. Table 15 shows that the full model under equal weights reached 40.7%, which is lower than the tuned configuration (43.8%). Importantly, the incremental trend remains consistent across both setups: $\mathcal{L}^{t}$, $\mathcal{L}_{aw}$, and $\mathcal{L}_{agn}$ each contribute positively to performance. This confirms that the effectiveness of MuSACo does not hinge on carefully tuned weights, but rather that appropriate weighting further amplifies the gains. Together, these results underline the robustness and general applicability of the proposed loss formulation.
10.7 Target Pseudo-label Progression During Training
We evaluate the reliability of co-training by tracking pseudo-label (PL) accuracy during training. Results are reported for two challenging StressID subjects: ctzy (Sub-4) and x1q3 (Sub-10), which exhibit high label imbalance. To stabilize PL quality, we adopt a confidence-annealing schedule, initializing $\tau_{pl}=0.95$ and gradually reducing it by 0.01 every $N$ epochs (e.g., 0.94 at Ep-4, 0.93 at Ep-8). As shown in Fig. 7, PL accuracy steadily improves across epochs, demonstrating the effectiveness of iterative refinement under this schedule. For Sub-4, PL accuracy improves from 0.976 at initialization to 0.996 after 8 epochs, while for Sub-10, which is highly imbalanced, accuracy increases from 0.602 to 0.666. These trends indicate that co-training with confidence annealing reliably enhances PL quality, even under challenging conditions.
Figure 7: Target pseudo-labels accuracy during training.
Settings Methods Target Subjects
Sub-1 Sub-2 Sub-3 Sub-4 Sub-5 Sub-6 Sub-7 Sub-8 Sub-9 Sub-10 Avg Acc
Lower Bound
Visual-only
Physio-only
Fusion
37.8
41.7
36.0
27.7
30.1
28.0
23.3
30.5
26.0
22.6
33.3
35.0
27.5
36.7
35.0
21.6
30.9
15.0
37.0
34.7
43.0
20.0
38.0
40.0
22.0
37.3
35.0
24.0
28.0
28.0
26.4
34.1
32.1
$M$-UDA
MMD [sejdinovic2013equivalence]
MuSACo (UDA)
40.3
39.0
31.2
35.0
32.3
34.0
26.3
32.0
33.3
29.0
24.1
19.0
34.3
38.9
39.5
37.0
37.3
31.0
20.0
29.9
31.9
32.5
$M$-MSDA
Sub-basedtop-k [zeeshan2024subject]
MuSACo (MSDA)
45.0
42.0
35.0
38.0
39.0
51.0
37.0
30.0
34.0
28.0
32.0
39.0
42.0
43.0
36.0
35.0
27.0
43.0
30.0
33.0
35.7
38.2
Table 19: Subject-wise accuracy of the BioVid dataset using the ViT backbone across 10 target subjects.
Settings Methods Target Subjects
Sub-1 Sub-2 Sub-3 Sub-4 Sub-5 Sub-6 Sub-7 Sub-8 Sub-9 Sub-10 Avg F1
Lower Bound
Visual-only
Physio-only
Fusion
35.6
31.0
23.9
27.5
21.9
27.1
22.5
23.1
22.8
12.5
24.5
22.6
19.7
27.9
33.5
19.6
23.6
9.5
35.3
23.9
32.9
15.6
27.3
34.6
14.9
28.2
29.1
23.2
19.1
26.1
22.6
25.0
26.2
$M$-UDA
MMD [sejdinovic2013equivalence]
MuSACo (UDA)
31.7
33.6
33.1
34.6
28.6
25.9
20.5
26.9
33.8
25.9
23.5
18.4
37.7
38.4
35.6
33.8
29.3
31.3
18.1
29.8
29.2
29.9
$M$-MSDA
Sub-basedtop-k [zeeshan2024subject]
MuSACo (MSDA)
36.9
33.9
27.2
27.6
34.6
38.7
32.9
26.4
20.6
24.9
27.4
33.9
34.5
41.5
28.0
30.1
13.1
37.6
22.6
29.8
27.8
32.4
Table 20: Subject-wise F1 scores of the BioVid dataset using the ViT backbone across 10 target subjects.
Settings Methods Target Subjects
Sub-1 Sub-2 Sub-3 Sub-4 Sub-5 Sub-6 Sub-7 Sub-8 Sub-9 Sub-10 Avg Acc
Lower Bound
Visual-only
Physio-only
Fusion
57.0
57.8
48.3
18.3
84.1
14.0
41.7
60.7
49.3
68.3
28.8
78.3
62.4
60.1
66.0
47.5
69.0
49.2
84.2
84.1
74.0
47.0
51.2
51.3
65.5
34.4
64.2
69.7
29.4
73.5
56.2
56.0
56.8
$M$-UDA
MMD [sejdinovic2013equivalence]
MuSACo (UDA)
53.0
64.0
23.0
20.0
36.4
38.0
82.0
79.0
69.0
65.0
47.6
50.0
78.0
82.0
46.6
50.0
77.0
66.0
66.0
68.0
57.9
58.2
$M$-MSDA
Sub-basedtop-k [zeeshan2024subject]
MuSACo (MSDA)
58.0
38.0
15.0
38.0
50.0
87.0
72.0
100.0
62.0
67.0
82.0
67.0
94.0
85.0
82.0
82.0
66.0
66.0
67.0
68.0
64.8
69.8
Table 21: Subject-wise accuracy of the StressID dataset across 10 target subjects using the ViT backbone.
Settings Methods Target Subjects
Sub-1 Sub-2 Sub-3 Sub-4 Sub-5 Sub-6 Sub-7 Sub-8 Sub-9 Sub-10 Avg F1
Lower Bound
Visual-only
Physio-only
Fusion
55.2
36.6
47.0
17.3
45.7
13.0
32.5
37.8
36.4
67.9
22.4
77.4
46.1
37.5
53.4
46.3
40.8
47.1
76.4
45.7
64.2
32.9
33.9
42.8
39.6
25.6
39.6
41.1
22.7
42.3
45.5
34.9
46.3
$M$-UDA
MMD [sejdinovic2013equivalence]
MuSACo (UDA)
51.6
62.6
22.6
19.3
27.5
28.5
72.6
77.4
55.2
49.3
42.1
48.5
72.5
72.8
44.6
34.8
44.0
39.5
39.5
40.2
47.2
47.3
$M$-MSDA
Sub-basedtop-k [zeeshan2024subject]
MuSACo (MSDA)
56.0
28.5
13.8
36.4
46.5
85.2
71.0
98.8
38.0
66.3
80.1
66.2
87.3
45.7
81.2
81.2
39.5
39.5
39.9
40.2
55.3
58.8
Table 22: Subject-wise F1 scores of the StressID dataset across 10 target subjects using the ViT backbone.
11 Subject-wise F1 Scores
11.1 Stress-ID
Table 16 presents subject-wise F1 scores on the StressID dataset. While the dataset has a binary label distribution (stress vs. no-stress), clear variability emerges across subjects. For example, lower-bound fusion consistently yields higher F1 than either modality alone, showing the complementary nature of visual and physiological signals. However, without adaptation, predictions remain unbalanced across subjects, with several experiencing sharp drops (e.g., Sub-2, Sub-4). UDA methods improve F1 moderately on some subjects, but results remain inconsistent. In contrast, MuSACo (MSDA) achieves the best overall performance, boosting recall on harder subjects (e.g., Sub-9) and producing the highest average F1 across all targets. These findings underscore the importance of F1 as a complementary metric to accuracy, particularly for stress recognition, where subject-specific imbalance strongly affects model robustness.
11.2 Biovid
Tab. 17 reports subject-wise F1 scores on the BioVid dataset. Our analysis shows that MuSACo particularly struggles on subjects where the Physio-only baseline itself is weak (e.g., Sub-4, Sub-5, Sub-7). This suggests that physiological variability across individuals is the key limiting factor: when physio cues are unreliable, visual signals alone are too subtle to compensate, and fusion inherits this weakness. These subject-specific outcomes highlight an inherent challenge of BioVid, where the modality that is usually strongest (physio) can be inconsistent for some individuals, constraining overall adaptation performance. Nevertheless, MuSACo achieves consistent performance across all subjects and outperforms competing methods on 6 out of 10 targets, leading to the best overall F1 score.
11.3 BAH
Tab. 18 reports subject-wise F1 scores on the BAH dataset. While the visual modality is generally stronger than audio, naive fusion provides slight gains by leveraging complementary information across modalities. MuSACo (MSDA) achieves the best overall performance, consistently surpassing both lower-bound baselines and existing UDA/MSDA approaches, demonstrating the effectiveness of source subject selection and multimodal adaptation in more challenging, in-the-wild conditions.
12 Evaluation on Additional Backbone
In the main paper, all experiments are conducted using a ResNet-18 backbone to ensure fair comparison with prior work and maintain a computationally efficient setup. However, MuSACo is inherently backbone-agnostic and does not rely on architectural properties specific to convolutional networks. To verify that MuSACo extends beyond ResNet-based architectures, we additionally evaluate the method using a Vision Transformer (ViT). We use the same MuSACo protocol as before for all the experiments, including the same setup for baseline methods. Biovid. Tables 19 and 20 report the subject-wise Accuracy and F1 scores across 10 target subjects. MuSACo consistently surpasses the baseline UDA and MSDA methods, achieving the highest average Accuracy and F1. Notably, MuSACo maintains strong performance even in subjects where the baselines degrade, demonstrating stable target adaptation. StressID. Tables 21 and 22 present the subject-wise Accuracy and F1 results. MuSACo consistently outperforms the baseline UDA and MSDA methods across most subjects, achieving the highest average Accuracy and F1. In particular, MuSACo demonstrates strong improvements on challenging subjects such as Sub-3, Sub-4, and Sub-6, where existing approaches show significant performance drops. These results further validate that the effectiveness of MuSACo is preserved even when replacing ResNet-18 with a transformer-based backbone, confirming that the proposed method is robust and generalizes well across different architectural families.
13 MuSACo